Three charges are at the corners of an isosceles triangle. The ±5.00µC charges from a dipole that is 3 cm apart.

  1. Find the force (magnitude and direction) the -10.00µC charge, 2 cm from the dipole, exerts on the dipole.
  2. For an axis perpendicular to the line connecting the ±5.00µC charges at the mid-point of this line, find the torque (magnitude and direction) exerted on the dipole by the 10.00µC charge.
1
Expert's answer
2021-02-15T17:38:41-0500

(1) Electric field due to dipole at equatorial point is given by, "E = \frac{p}{4\pi \epsilon_0 (r^2+a^2)^({3}/{2})}"

where p is the dipole moment, r is the separation of the charge from one charge of the dipole, 'a' is the semi length of the dipole.


Force is given by, "F = qE = \frac{qp}{4\pi \epsilon_0 (r^2+a^2)^({3}/{2})}"

Putting values we get,

"F = qE = \frac{qp}{4\pi \epsilon_0 (r^2+a^2)^({3}/{2})} = \:\frac{9\cdot 10^9\cdot \left(5\cdot 10^{-5}\cdot 0.03\right)\cdot 10\cdot 10^{-6}}{\:\left(0.02^2+0.03^2\right)^{\left(3/2\right)}} = 2880.17 N" towards negative charge of the dipole parallel to the dipole.


(2) Torque is given by, "\tau = rFsin\theta=pEsin\theta" where E is the external electric field

and "rsin\theta" is the perpendicular separation of the two forces.

Here Both forces are intersecting at charge "10\mu C" . So torque will be zero.